Consider the two vectors, A = 18i + 39j B = 7.7i - 21.5j

What is the magnitude of the vector 3A - B? What is the direction of the vector 3A - B?
Answer #1

3A - B = 3*[18i + 39j] - [ 7.7i - 21.5j ]

= 54i + 117j - 7.7i + 21.5j

   = 46.3i + 138.5 j

magnitude = sqrt(46.3^2+138.5^2) = 146

direction = arctan(138.5/46.3) = 71.5 degree with x axis